Ask
your own question, for FREE!
Mathematics
18 Online
If you roll a fair die twice, what is the probability of getting an even number at least once? [will medal]
Still Need Help?
Join the QuestionCove community and study together with friends!
how many even numbers are there on a dice
tell me @tishany
well there are three 2,4,6 and there are three odds 1,3,5 which means you have a 50 50 chance
@tishany Cody is helping you, could you please respond?
my laptop froze and the test closed
Still Need Help?
Join the QuestionCove community and study together with friends!
@cody77445
its fine math mate stuff like that happens to me all the time
Hint for your future problems. When you deal with probabilities of "at least once", you can replace P(at least once) = 1 - P(not even once).
thank you @cody77445 @mathmate
@tishany Just sent you a hint for future problems. You could finish the current problem anyway for your self, which is 1-(1/2)*(1/2)
Still Need Help?
Join the QuestionCove community and study together with friends!
np
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
12 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal